OpenStudy (anonymous):

Activity title: solving problems for uniformly accelerated rectilinear motion learning target: to calculate displacement,velocity,acceleration and time using the four kinematic equations. 1. A french high-velocity train accelerated uniformly from 250km/hr to 300 km/hr as it sped through a distance of 50 km.finds its acceleration.

OpenStudy (anonymous):

You know the initial speed, and you know the final speed, as well as teh distance travelled. Thus you will want to use the equation of motion \[v^2=u^2+2as\]where \(v\) is the final velocity, \(u\) is the initial velocity, \(s\) is the displacement (or distance travelled) and \(a\) is the acceleration. The fact that the train is French should not effect the calculation :P

OpenStudy (anonymous):

can u give me the example of this answer im slow

OpenStudy (anonymous):

300^2 = 250^2 +2(50)a , solve for a and you will have you answer in km/hr^2

OpenStudy (anonymous):

Just as cmcinnnis says. v = 300, u = 250, s = 50, and acceleration is given by \(a\)

OpenStudy (anonymous):

jonny answered right

OpenStudy (anonymous):

so the final asnwer is 0.02 m/s ?

OpenStudy (anonymous):

we need conversion